There are 100g of sugar dissolved in 1.5l of water. How much sugar is present if you take 0.5qt of the liquid from the solution?
Answer by Theo(13342)

you have 100 grams of sugar dissolved in 1.5 liters of water.
you want to remove 1 quart of the solution.

you have 100 grams of sugar dissolved in 1.5 liters of water.

the ratio of grams of sugar to liters of water is 100 / 1.5.

1 quart is equal to 0.946353 liters.

1/2 of a quart is therefore equal to 1/2 * .946353 liters which is equal to .4731765 liters.

when you remove 1/2 of a quart of liquid, you are also removing 100 / 1.5 * .4731765 = 31.5451 grams of sugar.

this assumes the ratio of sugar to liquid remains the same.

the amount of liquid that is left is 1.5 - .4731765 = 1.0268235 liters.

the amount of sugar that is left is 100 / 1.5 * 1.0268235 = 68.4549 grams of sugar.

add up the sugar that is removed and the sugar that is left and you get 31.5451 + 68.4549 = 100 grams of sugar.

add up the liquid that is removed and the liquid that is left and you get .4731765 + 1.0268235 = 1.5 liters of water.

the assumption is that, when you remove the water, you are also removing the sugar that is dissolved in the water.

the question is: How much sugar is present if you take 0.5qt of the liquid from the solution?

if you are asking how much sugar is present in the half a quart that is removed, then the answer is 31.5451 grams.

if you are asking how much sugar is left in the solution after you have removed half a qoart, then the answer is 68.4549 grams.

hopefully, i understood your problem correctly.

let me know if there's another interpretation that i might have missed.